# A flexible, GPU-accelerated approach for the joint characterization of LISA instrumental noise and Stochastic Gravitational Wave Backgrounds

### v1

@@ I Introduction
In addition, LISA will be sensitive to stochastic gravitational-wave backgrounds (SGWBs) of astrophysical or cosmological origin. Astrophysical stochastic backgrounds are generated by the incoherent superposition of GW signals from a large number of individual sources. Astrophysical populations that could produce detectable stochastic backgrounds for LISA include Galactic [16] and extragalactic [17] WDBs, SBHBs [18] and EMRIs [19, 20, 21]. The WDB background in particular is expected to dominate over the expected instrumental noise [22] between $\sim 0.5$ and $\sim 3$ mHz. SGWBs in the LISA band could also be of cosmological origin [23]. Possible cosmological sources include First Order Phase Transitions (FOPTs), which are expected in many theories that extend physics beyond the Standard Model [24, 25, 26], Cosmic Strings [27], primordial fluctuations in the early universe [28], and Primordial Black Holes (PBHs) [29, 30].
@@
In this work, we focus on the task of simultaneously characterizing the instrumental noise and possible SGWBs present in the data. Previous studies have tackled this problem by using relatively simple parametrized templates with $2-12$ parameters to describe the spectral shape of the instrumental noise and of the different SGWBs considered (see, e.g., Refs. [35, 36, 37]). Nevertheless, we expect that the instrumental noise entering the LISA data streams will be significantly more complex and hard to fully model analytically. The LISA-Pathfinder (LPF) mission [38, 39] has already shown a similar behaviour, with a significant fraction of the total measured noise that could not be explained by on-board measurements of individual noise components. As the goal of LPF was to test the behavior of (part of) the key LISA hardware, we can already expect the same to happen for LISA, which is a significantly more complex system than LPF. Therefore, we must be prepared to account for uncertainties in the noise models we will ultimately use in data analysis pipelines, which will come at the cost of affecting the detection and characterisation of SGWBs [40]. Moreover, the different SGWBs described are expected to have different spectral shapes. Fitting them would require the inclusion of models for all of them in our pipelines, which is not only impractical but would also require a complete SGWB template bank. Since our knowledge of the exact spectral shape of the different SGWBs is limited, we cannot even compile a complete catalogue of all the possible sources and associated spectral shapes. For this reason, robust and flexible signal reconstruction algorithms are of pivotal importance to successfully extract the information carried by these persistent signals.
@@ II Gravitational Wave signals and noise models
In this section, we report the properties of the instrumental noise and SGWBs considered in this work. We then combine them to describe the datasets simulated for the analysis.
@@ II.1 Time-delay Interferometry
While many different TDI combinations can be (and have been) constructed [51, 68, 64], the Michelson variables $X$, $Y$, and $Z$ are widely employed in the community [69, 70, 71, 72]. These can be further combined to form the three noise-orthogonal $A,\,E,\,T$ channels [62, 73]. In principle, this TDI basis does not present cross-correlations between the different channels, significantly simplifying GW likelihoods and speeding up their evaluation. However, this orthogonality holds only under the assumptions of equal arm lengths, uncorrelated noise contributions, and identical OMS and TM noises in all the links. This TDI basis is constructed as a linear combination of the Michelson channels:
@@
[…]
The other two TDI variables $Y_{2}$ and $Z_{2}$ are defined in a similar way to $X_{2}$ by cyclic permutations of the three satellites. In the above, $D_{ij}$ is an operator that applies a constant delay equal to $\tau_{ij}=L_{ij}/c$, with $\tau_{ij}$ being the light travel time between spacecrafts $j$ (the emitter) and $i$ (the receiver):
@@
In the frequency domain, its effect can be expressed as a multiplication by the phase factor $\mathcal{F}(D_{ij})=e^{-i2\pi f\tau_{ij}}$. The quantities $\eta_{ij}(t)$ represent the single-link measurements between spacecrafts [64], describing the relative phase difference between the laser beam emitted from spacecraft $j$ at time $t-\tau_{ij}$ and the laser local to spacecraft $i$ at time $t$. In this work, we assume equal and constant arm lengths, so that $L_{ij}=L=2.5\times 10^{9}\,\rm m$ for all the $i,j$ indices.
II.2 Noise model
@@
We assume the Scirdv1 noise curve [74, 22], such that the TM acceleration component is described by:
$\expectationvalue{\tilde{x}^{g}_{ij}(f)\tilde{x}^{g\,\dagger}_{lm}(f^{\prime})}=\frac{1}{2}\delta_{il}\delta_{jm}\delta(f-f^{\prime})S_{\text{TM}}(f)$ (4)
@@
×$\qquad\qquad\quad\times\left(\frac{1}{2\pi fc}\right)^{2}\!.$
We set the test mass noise amplitude to $A_{\rm TM}=2.4\times 10^{-15}\,\rm m\,s^{-2}\,Hz^{-1/2}$. Conversely, the OMS component is:
$\expectationvalue{\tilde{x}^{m}_{ij}(f)\tilde{x}^{m\,\dagger}_{lm}(f^{\prime})}$ $=\frac{1}{2}\delta_{il}\delta_{jm}\delta(f-f^{\prime})S_{{\rm OMS}}(f)$ (5)
@@
Through this work, we use $A_{\rm OMS}=7.9\times 10^{-12}\,\rm m\,Hz^{-1/2}$. Daggers represent complex conjugates, and we use the notation $\expectationvalue{\cdot}$ to indicate expectation values. $S_{\rm TM}$ and $S_{\rm OMS}$ represent the one-sided Power Spectral Densities (PSDs) of the two components. We assume that the PSDs of all six test mass noise terms and, separately, all six OMS noise terms are the same, and the noise components are independent. These assumptions mean that the noises in the $A$, $E$ and $T$ channels will be uncorrelated.
@@
II.3 SGWB model
Stochastic backgrounds in the LISA band can have a large variety of spectral shapes, depending on the physical processes that generate them [75]. For the purposes of this work, we only consider SGWBs whose energy density can be described by a power-law:
@@ II.4 Data generation
We test our pipeline on two different setups to study its performance under different conditions. In the first case, we generate and analyze data consistent with a stochastic background from cosmic strings just outside the reach of SKA [81], choosing $G\mu=10^{-13}$. We inject the signal in a realization of LISA instrumental noise generated according to the LISA PSD introduced in Eq. (6). Conversely, in the second case, we analyze a noise realization that is produced from a PSD whose spectral shape differs from what we expect from the instrument. This is described in Sec. IV.2.
@@ III.1 Power spectral density breakdown
We perform our analysis using multiple channels. Thus, Eq. (11) applies to each of them. We can write the data vector of the Fourier transformed TDI variables as $\boldsymbol{\tilde{\rm d}}\equiv(\tilde{A},\,\tilde{E},\,\tilde{T})^{T}$ and then, at each frequency, we can define the $3\times 3$ periodogram matrix $\boldsymbol{\rm P}(f)\equiv\boldsymbol{\tilde{\rm d}}(f)\boldsymbol{\tilde{\rm d}}(f)^{\dagger}$. Following [43], we include in the Fourier variables the normalization factor $\sqrt{2/(N\,f_{s})}$, with $N=T_{\rm obs}f_{s}$ the length of the time series. In this way, the periodogram is directly comparable to the one-sided PSD.
@@
Here $S_{n,ii}(f)$ and $S_{h,ii}(f)$ represent respectively the noise and background contribution to the overall one-sided PSD in each channel. We account for the non-perfect knowledge of their spectral shape, expressing them as
@@
For each component, we have factored the overall PSD as the product of a parametrized template $S_{n/h}^{\rm template}(f|\boldsymbol{\lambda}_{n/h})$ and a perturbation $\mathcal{P}_{n/h}(f|\boldsymbol{\gamma}_{n/h})$. We perform our inference algorithm on the parameters of both the templates and the perturbations, $\boldsymbol{\theta}=\{\boldsymbol{\lambda},\,\boldsymbol{\gamma}\}=\{\boldsymbol{\lambda}_{n},\,\boldsymbol{\lambda}_{h},\,\boldsymbol{\gamma}_{n,ii},\,\boldsymbol{\gamma}_{h}\}$. This factorization results in two correlated terms, but we choose it to allow the templates to catch the overall frequency evolution of the PSD and let eventual deviations from it be absorbed into $\mathcal{P}_{n/h}(f)$. In this work, we assume three independent perturbations for the noise contributions across the three TDI channels, whereas the background spectral shape is perturbed in a consistent (i.e., shared) way. Hence, the presence of the indices ${ii}$ only in the exponent of Eq. (14a). This assumption, which translates into using four different perturbations, is based on the hypothesis that the transfer function of the signal will be well-known. While the same holds for any noise source entering the measurement chain, we cannot know a-priori the full set of noises that will be present in the data stream. In addition, we will not have access to all the physical parameters that constitute the noises we expect to be there (see, e.g., Refs. [1] and [84]). Thus, we decided to be fully generic only in the noise contribution to the overall PSD in each channel. This approach has a caveat: the $AET$ basis is noise-orthogonal only under the assumption of a perfect equilateral constellation where all the secondary noise sources are exactly equal in the three spacecrafts [85], thus using independent splines in the three channels may seem a violation of this assumption. For a complete analysis, we should also include perturbations in the off-diagonal components of the noise matrix, which would increase the number of perturbations for the noise contributions by another three terms. These would also account for cross-correlations between the noise and SGWB components at high frequencies [40], but the amplitude of the SGWB considered in our analysis is significant only around a few millihertz, where the cross-correlations are a factor of $10^{-2}-10^{-3}$ smaller than the diagonal terms. We can therefore safely disregard off-diagonal terms. For a full analysis in the Michelson TDI basis, the (hermitian) covariance matrix $\mathbf{C}_{XYZ}(f)$ would be dense with $6$ independent components, implying the evaluation of $6$ splines at every likelihood call only for the noise contribution, significantly increasing the computational cost. Therefore, we work in this framework, ignoring the off-diagonal terms, for computational simplicity. Having generated the frequency domain data directly in the $AET$ basis, we know the noise-orthogonal assumption holds. We leave a full covariance treatment of the three noise channels for future analysis. We use as perturbations $\mathcal{P}_{n/h}(f)$ realizations of Akima splines [86], as they tend to show a smoother behavior with respect to Cubic splines. To be as generic as possible, we do not fix the number or the position of the spline knots beforehand. Instead, we use a Reversible Jump Markov Chain Montecarlo (RJ-MCMC) scheme [87] to dynamically infer the most probable number of knots. We use the parallel tempered ensemble sampler Eryn [88, 89, 90] for this. Our knot placement scheme can be described as follows:
@@
We set the minimum and maximum number of in-between knots to 00 and $10$, respectively. Since our Akima splines implementation uses boundary conditions that require at least $4$ points to be computed, we rely on a linear interpolant for cases with less than two in-between knots.
@@ III.3 MCMC setup
We run Eryn using 20 different temperatures with 20 walkers each. We initialise our walkers, drawing their position in the parameter space randomly from the prior. The only exception is for the spline knots, whose initial amplitudes are drawn from a multivariate normal centred on zero $\mathcal{N}(\boldsymbol{\mu}=\mathbf{0},\,\boldsymbol{\Sigma}=\boldsymbol{\rm I}\times 10^{-8})$. This choice follows our interpretation of the splines as “small” perturbations of parametrised templates. For each of our parameter estimation (PE) runs, we monitor the convergence of the samples by making sure that the chains are longer than $200\times\max(\hat{\tau})$, with $\hat{\tau}$ being the average autocorrelation time across all the model parameters [95]. We discard (at least) the first $5\times\max(\hat{\tau})$ samples to account for the burn-in phase, and we thin the chains by a factor of $0.5\times\min(\hat{\tau})$.
@@
where $d$ represents the observed data, $\boldsymbol{\theta}$ the model parameters, $\pi(\boldsymbol{\theta})$ their prior and $\mathcal{L}$ the assumed likelihood. We report our prior distributions in Tab. 1. The evidence is usually expensive to compute, and being a normalization constant, it can safely be disregarded in parameter estimation studies. However, it is required to perform Bayesian model selection. The ratio of the evidence computed for two competing models, $M_{1}$ and $M_{2}$, is called the Bayes factor, $\mathcal{B}_{12}=Z_{1}/Z_{2}$. A large Bayes factor indicates that model $M_{1}$ describes the observed data better than $M_{2}$. Examples of the use of Bayes factor’s for model selection in the field of GW astronomy can be found in Refs. [97, 98, 99]. Following [96], the stepping-stone approximation for the evidence can be expressed as:
$\log{\hat{Z}}=$ $\sum_{k=1}^{K}\log\sum_{i=1}^{n}\mathcal{L}(d\,|\,\boldsymbol{\theta}^{i}_{\beta_{\,k-1}},\,M)^{\beta_{k}-\beta_{\,k-1}}$ (16)
@@
To further reduce the computational cost of our runs, we average the periodograms over $N_{\rm bins}$ consecutive and non-overlapping frequency bins, and we sample using a complex Wishart likelihood, as described in Ref. [43]. We define the averaged periodogram in a segment $i$ as
@@
where $n_{i}$ is the number of frequency samples in a bin centred on $f_{i}^{\rm bin}$. Introducing the number of Degrees of Freedom (DoFs) in each bin $\nu(f_{i}^{\rm bin})=n_{i}$, the matrix $\boldsymbol{\rm Y}(f)=\nu(f)\boldsymbol{\rm P}(f)$ would follow the aforementioned complex Wishart distribution. When working with time domain data, we apply a window function to the timeseries before converting it into the frequency domain. This would introduce correlations between frequency bins, effectively reducing the number of DoFs $\nu(f)$. Since we generate our data directly in the frequency domain, this step is not necessary. Keeping only the terms that depend on the model parameters, for each bin we can write:
@@
where $\rm{tr}(\cdot)$ denotes the trace operator, and ||$\cdot$|| ⋅ | represents the determinant of a matrix. Combining all the bins together, our likelihood reads out:
@@
For completeness, we include in our code the possibility of using directly the totality of the frequency domain data and the Whittle likelihood [100]. This represents the standard in the context of GW astronomy, and it has already been widely implemented in the literature (see, e.g., Ref. [44], and Ref. [101] for a more comprehensive SGWB likelihoods review). We have checked that the two methods yield consistent results.
The averaging procedure, combined with hardware acceleration described in Sec. III.2, allows our runs to converge and accumulate enough samples within a day on an NVIDIA A100-SXM4 GPU despite the large number of parallel walkers and the dimensionality of the parameter space.
Table 1: Prior distributions used for the MCMC runs presented in this work. $A$ and $n$ represent the amplitude and slope of the power-law in Eq. (7). We use uniform distributions for most of the sampling parameters. The prior distribution assumed for the weight of each spline knot and edge is a normal distribution centered on $\mu=0$. In each channel and for each contribution, we assume the same priors on $f_{k}$ and $w_{k}$.

### v2

@@ I Introduction
In addition, LISA will be sensitive to stochastic gravitational-wave backgrounds (SGWBs) of astrophysical or cosmological origin. Astrophysical stochastic backgrounds are generated by the incoherent superposition of GW signals from a large number of individual sources. Astrophysical populations that could produce detectable stochastic backgrounds for LISA include Galactic [16] and extragalactic [17] WDBs, SBHBs [18] and EMRIs [19, 20, 21]. The WDB background in particular is expected to dominate over the expected instrumental noise [22] between $\sim 0.5$ and $\sim 3$ mHz. SGWBs in the LISA band could also be of cosmological origin [23]. Possible cosmological sources include First Order Phase Transitions (FOPTs), which are expected in many theories that extend physics beyond the Standard Model [24, 25, 26], Cosmic Strings [27], primordial fluctuations in the early universe [28], and Primordial Black Holes (PBHs) [29, 30, 31].
@@
In this work, we focus on the task of simultaneously characterizing the instrumental noise and possible SGWBs present in the data. Previous studies have tackled this problem by using relatively simple parametrized templates with 2122-12 parameters to describe the spectral shape of the instrumental noise and of the different SGWBs considered (see, e.g., Refs. [36, 37, 38]). Past works also include a templated analysis of the WDB background anisotropy [39]. Nevertheless, we expect that the instrumental noise entering the LISA data streams will be significantly more complex and hard to fully model analytically. The LISA-Pathfinder (LPF) mission [40, 41] has already shown a similar behaviour, with a significant fraction of the total measured noise that could not be explained by on-board measurements of individual noise components. As the goal of LPF was to test the behavior of (part of) the key LISA hardware, we can already expect the same to happen for LISA, which is a significantly more complex system than LPF. Therefore, we must be prepared to account for uncertainties in the noise models we will ultimately use in data analysis pipelines, which will come at the cost of affecting the detection and characterisation of SGWBs [42]. Moreover, the different SGWBs described are expected to have different spectral shapes. Fitting them would require the inclusion of models for all of them in our pipelines, which is not only impractical but would also require a complete SGWB template bank. Since our knowledge of the exact spectral shape of the different SGWBs is limited, we cannot compile a complete catalogue of all the possible sources and associated spectral shapes. For this reason, robust and flexible signal reconstruction algorithms are of pivotal importance to successfully extract the information carried by these persistent signals.
@@ II Gravitational Wave signals and noise models
In this section, we report the properties of the instrumental noise and SGWBs considered in this work. We then combine them to describe the datasets simulated for the analysis. In particular, we describe in Secs. II.2 and II.3 the properties of the parametrized noise and SGWB templates, respectively, that we use as baselines in our flexible model.
@@ II.1 Time-delay Interferometry
While many different TDI combinations can be (and have been) constructed [53, 69, 65], the Michelson variables $X$, $Y$, and $Z$ are widely employed in the community [70, 71, 72, 73]. These can be further combined to form the three noise-orthogonal A,E,TA,\,E,\,T channels [63, 74]. In principle, this TDI basis does not present cross-correlations between the different channels, significantly simplifying GW likelihoods and speeding up their evaluation. However, this orthogonality holds only under the assumptions of equal arm lengths, uncorrelated noise contributions, and identical OMS and TM noises in all the links. This TDI basis is constructed as a linear combination of the Michelson channels:
@@
$\quad+D_{12}D_{21}D_{13}D_{31}D_{13}D_{31})\times(\eta_{12}+D_{12}\eta_{\,21}).$
In the above, the quantities $\eta_{ij}(t)$ represent the single-link measurements between spacecrafts [65], describing the relative phase difference between the laser beam emitted from spacecraft $j$ at time $t-\tau_{ij}$ and the laser local to spacecraft $i$ at time $t$. Here, $\tau_{ij}=L_{ij}/c$ represents the light travel time between spacecrafts $j$ (the emitter) and $i$ (the receiver). $D_{ij}$ is, instead, an operator that applies a constant delay equal to $\tau_{ij}$, with:
@@
In the frequency domain, its effect can be expressed as a multiplication by the phase factor $\mathcal{F}(D_{ij})=e^{-i2\pi f\tau_{ij}}$. In this work, we assume equal and constant arm lengths, so that $L_{ij}=L=2.5\times 10^{9}\,\rm m$ for all the $i,j$ indices. The other two TDI variables, $Y_{2}$ and $Z_{2}$, are defined similarly to $X_{2}$ by cyclic permutations of the three satellites.
II.2 Parametrized noise model
@@
In this work, we assume a noise curve [70, 75] such that the TM acceleration component is described by:
$\expectationvalue{\tilde{x}^{g}_{ij}(f)\tilde{x}^{g\,*}_{lm}(f^{\prime})}=\frac{1}{2}\delta_{il}\delta_{jm}\delta(f-f^{\prime})S_{\text{TM}}(f)$ (4)
@@
$\qquad\qquad\quad\times\left(\frac{1}{2\pi fc}\right)^{2}\!.$
We set the test mass noise amplitude to $A_{\rm TM}=2.4\times 10^{-15}\,\rm m\,s^{-2}\,Hz^{-1/2}$ [70]. Conversely, the OMS component reads:
$\expectationvalue{\tilde{x}^{m}_{ij}(f)\tilde{x}^{m\,*}_{lm}(f^{\prime})}$ $=\frac{1}{2}\delta_{il}\delta_{jm}\delta(f-f^{\prime})S_{{\rm OMS}}(f)$ (5)
@@
[…]
@@
II.3 Parametrized SGWB model
Stochastic backgrounds in the LISA band can have a large variety of spectral shapes, depending on the physical processes that generate them [76]. For the purposes of this work, we only consider SGWB templates whose energy density can be described by a power-law:
@@ II.4 Data generation
We test our pipeline on two different setups to study its performance under different conditions. In the first case, we generate data consistent with a parametrized stochastic background from cosmic strings just outside the reach of SKA [82], choosing $G\mu=10^{-13}$ in Eq. (8). We inject the signal in a realization of LISA instrumental noise generated according to the parametrized LISA PSD introduced in Eq. (6). Conversely, in the second case, we analyze a noise realization that is produced from a PSD whose spectral shape differs from what we expect from the instrument. This is described in Sec. IV.2.
@@ III.1 Power spectral density breakdown
We perform our analysis using multiple channels. Thus, Eq. (11) applies to each of them. We can write the data vector of the Fourier transformed TDI variables as $\boldsymbol{\tilde{\rm d}}\equiv(\tilde{A},\,\tilde{E},\,\tilde{T})^{T}$ and then, at each frequency, we can define the $3\times 3$ periodogram matrix $\boldsymbol{\rm P}(f)\equiv\boldsymbol{\tilde{\rm d}}(f)\boldsymbol{\tilde{\rm d}}(f)^{\dagger}$. Daggers represent conjugate transposes. Following [45], we include in the Fourier variables the normalization factor $\sqrt{2/(N\,f_{s})}$, with $N=T_{\rm obs}f_{s}$ the length of the time series. In this way, the periodogram is directly comparable to the one-sided PSD.
@@
Here $S_{n,ii}(f)$ and $S_{h,ii}(f)$ represent respectively the noise and background contribution to the overall one-sided PSD in each channel. We account for the non-perfect knowledge of their spectral shape, factoring the overall PSD as the product of a parametrized template $S_{n/h}^{\rm template}(f|\boldsymbol{\lambda}_{n/h})$ and a perturbation $\mathcal{P}_{n/h}(f|\boldsymbol{\gamma}_{n/h})$. We perform our inference algorithm on the parameters of both the templates ($\boldsymbol{\lambda}=\{\boldsymbol{\lambda}_{n},\,\boldsymbol{\lambda}_{h}\}$) and the perturbations ($\boldsymbol{\gamma}=\{\boldsymbol{\gamma}_{n},\,\boldsymbol{\gamma}_{h}\}$). We represent the entire parameter space with $\boldsymbol{\theta}$: $\boldsymbol{\theta}=\{\boldsymbol{\theta}_{n},\,\boldsymbol{\theta}_{h}\}=\{\boldsymbol{\lambda},\,\boldsymbol{\gamma}\}=\{\boldsymbol{\lambda}_{n},\,\boldsymbol{\lambda}_{h},\,\boldsymbol{\gamma}_{n,ii},\,\boldsymbol{\gamma}_{h}\}$. Therefore, each component reads:
@@
This factorization results in two correlated terms, but we choose it to allow the templates to catch the overall frequency evolution of the PSD and let eventual deviations from it be absorbed into $\mathcal{P}_{n/h}(f)$. In this work, we assume three independent perturbations for the noise contributions across the three TDI channels, whereas the background spectral shape is perturbed in a consistent (i.e., shared) way. Hence, the presence of the indices ii{ii} only in the exponent of Eq. (14a). This assumption, which translates into using four different perturbations, is based on the hypothesis that the transfer function of the signal will be well-known. While the same holds for any noise source entering the measurement chain, we cannot know a-priori the full set of noises that will be present in the data stream. In addition, we will not have access to all the physical parameters that constitute the noises we expect to be there (see, e.g., Refs. [1] and [85]). Thus, we decided to be fully generic only in the noise contribution to the overall PSD in each channel. This approach has a caveat: the $AET$ basis is noise-orthogonal only under the assumption of a perfect equilateral constellation where all the secondary noise sources are exactly equal in the three spacecrafts [86], thus using independent splines in the three channels may seem a violation of this assumption. For a complete analysis, we should also include perturbations in the off-diagonal components of the noise matrix, which would increase the number of perturbations for the noise contributions by another three terms. The off-diagonal terms would also account for cross-correlations between the noise and SGWB components at high frequencies [38, 42]. Cross correlations can be as large as $\sim 10\,\%$ of the diagonal terms [38]. These can have an impact on the distinguishability of instrumental noise and SGWBs, but from the results of Ref. [42] we do not expect our conclusions to be heavily affected. For completeness, this should be validated against a full analysis in the Michelson TDI basis, where the (hermitian , positive-definite) covariance matrix $\mathbf{C}_{XYZ}(f)$ is dense with $6$ independent components, since it is possible to construct a positive definite matrix whose elements are functions of $6$ independent, flexible terms [42]. This requests the evaluation of $6$ splines at every likelihood call only for the noise contribution, significantly increasing the computational cost. Here, we neglect off-diagonal terms for computational simplicity. Having generated the frequency domain data directly in the $AET$ basis, we know the noise-orthogonal assumption holds. We leave a full covariance treatment of the three noise channels for future analysis.
We use as perturbations $\mathcal{P}_{n/h}(f)$ realizations of Akima splines [87], which although only being $C^{1}$ differentiable, were introduced to produce smoother fits to data in the presence of rapidly varying second derivatives with respect to other widely used $C^{2}$ Cubic splines implementations [88, 89, 90]. A description of the Akima splines algorithm can be found in App. A. To be as generic as possible, we do not fix the number or the position of the spline knots beforehand. Instead, we use a Reversible Jump Markov Chain Montecarlo (RJ-MCMC) scheme [91] to dynamically infer the most probable number of knots. We use the parallel tempered ensemble sampler Eryn [92, 93, 94] for this. Our knot placement scheme can be described as follows:
@@
We set the minimum and maximum number of in-between knots to 0 and $10$, respectively. Since our Akima splines implementation uses boundary conditions that require at least $4$ points to be computed, we rely on a linear interpolant for cases with less than two in-between knots.
@@ III.3 MCMC setup
[…]
@@
where $d$ represents the observed data, $\boldsymbol{\theta}$ the model parameters, $\pi(\boldsymbol{\theta})$ their prior and $\mathcal{L}$ the assumed likelihood. We report our prior distributions in Tab. 1. The evidence is usually expensive to compute, and being a normalization constant, it can safely be disregarded in parameter estimation studies. However, it is required to perform Bayesian model selection. The ratio of the evidence computed for two competing models, $M_{1}$ and $M_{2}$, is called the Bayes factor, $\mathcal{B}_{12}=Z_{1}/Z_{2}$. A large Bayes factor indicates that model $M_{1}$ describes the observed data better than $M_{2}$. Examples of the use of Bayes factor’s for model selection in the field of GW astronomy can be found in Refs. [101, 102, 103], while applications of the stepping-stone method in this field can be found in Refs. [104, 105]. Following [100], the stepping-stone approximation for the evidence can be expressed as:
[…]
@@
To further reduce the computational cost of our runs, we do not sample using the totality of the frequency domain data and the Whittle likelihood [106]. This represents the standard in the context of GW astronomy, and it has already been widely implemented in the literature (see, e.g., Ref. [46], and Ref. [107] for a more comprehensive SGWB likelihoods review). Here, we average the periodograms over $N_{\rm bins}$ consecutive and non-overlapping frequency bins, and we sample using a complex Wishart likelihood, as described in Ref. [45]. We define the averaged periodogram in a segment $i$ as
@@
where $n_{i}$ is the number of frequency samples in a bin centred on $f_{i}^{\rm bin}$. Introducing the number of Degrees of Freedom (DoFs) in each bin $\nu(f_{i}^{\rm bin})=n_{i}$, the matrix $\boldsymbol{\rm Y}(f)=\nu(f)\boldsymbol{\rm P}(f)$ would follow the aforementioned complex Wishart distribution if $\boldsymbol{\tilde{\rm d}}(f_{k})$ were uncorrelated between frequency bins. When working with time domain data, we apply a window function to the timeseries before converting it into the frequency domain. This would introduce correlations between frequency bins, effectively reducing the number of DoFs $\nu(f)$. Since we generate our data directly in the frequency domain, this step is not necessary. Keeping only the terms that depend on the model parameters, for each bin we can write:
@@
where $\rm{tr}(\cdot)$ denotes the trace operator, and $|\cdot|$ represents the determinant of a matrix. Combining all the bins together, our likelihood reads out:
@@
In the limit of $n_{i}=1$ for all $i$, having only one frequency sample per bin, the complex Wishart likelihood converges to the Whittle likelihood [45]. For completeness, we include in our code the possibility of using directly the latter with the totality of the frequency domain data. We have checked that the two methods yield consistent results.
Table 1: Prior distributions used for the MCMC runs presented in this work. AA and $n$ represent the amplitude and slope of the power-law in Eq. (7). We use uniform distributions for most of the sampling parameters. The prior distribution assumed for the weight of each spline knot and edge is a normal distribution centered on $\mu=0$. In each channel and for each contribution, we assume the same priors on $f_{k}$ and $w_{k}$.